OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 11 + 2 the square root of 11 22 the square root of 22 2 the square root of 22 3 the square root of 11

OpenStudy (anonymous):

am I doing each as a diffrent problem or all together as one?

OpenStudy (anonymous):

\[\sqrt{11}+2\sqrt{11}=3\sqrt{11}\] just like \[x+2x=3x\]

OpenStudy (anonymous):

all together

OpenStudy (anonymous):

where did you get the 3 from

OpenStudy (anonymous):

the rounded answer I get is six I know this is wrong so I can't help what class is this?

OpenStudy (anonymous):

this is algebra 1

OpenStudy (anonymous):

o thats the class im in too probably why I can't help u r u in connections?

OpenStudy (anonymous):

no i am not in connections well do you know anyone who can help me

OpenStudy (anonymous):

no sorry :(

OpenStudy (anonymous):

okay never mind i got it thank you anyway